OpenStudy (anonymous):

3^(sin2x+cos2x)+3^(3−sin2x−cos2x)=28 show that either cos x = 0 or tan x = –1

OpenStudy (anonymous):

3 ^(2sinxcosx + 2cos^2x - 1) = 3^-1) = 1/3 3^(3 - 0 - (-1)) = 3^4 = 81 Sorry sandee but I cant get 28 sin2x = 2sinxcosx and cos2x = 2 cos^2x - 1 and if soc x = 0 then first = 0 and second = 2(0) -1 = -1 maybe i've made a mistake - i have to go now